OpenStudy (anonymous):

Solve: 8b + 20 ≥ 5b – 10. A. b ≤ 10/3 B. b ≤ 10 C. b ≥ -10 D. b ≥ -10/3

OpenStudy (anonymous):

d

OpenStudy (whpalmer4):

D is incorrect. \[8b+20\ge 5b-10\]Subtract \(5b\) from each side: \[8b-5b+20\ge 5b-5b-10\]\[3b+20\ge -10\]Subtract 20 from each side:\[3b+20-20\ge-10-20\]\[3b\ge-30\]Now divide both sides by 3. What do you get?